OpenStudy (anonymous):

Walter took out a $6,000 loan for six years. He is being charged 6 percent interest, compounded annually. Calculate the total amount he will pay.

OpenStudy (anonymous):

compound interest formula = P * (1+ (r/100))^n p= principle, n= no of years, r = rate of interest substituting this we get CI = 6000 * (1+ (6/100))^6 solving this we ll get 8511.11 so he have to pay back $8511.11